Changes between Version 1 and Version 2 of Research/LhARA/RadiationBiology/Meetings/2025-05-01


Ignore:
Timestamp:
May 1, 2025, 2:41:24 PM (2 weeks ago)
Author:
longkr
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • Research/LhARA/RadiationBiology/Meetings/2025-05-01

    v1 v2  
    11= LhARA radiobiology; meeting: 10Apr25; 14:00 BST =
    22
    3 **ZOOM:** https://cern.zoom.us/j/61351195916?pwd=T0acw3Tid5bW7EQdGMmefUj2XIr7TY.1
     3**Present:** RG, CW, CH, RR, MH, RW, CD, KL, ND
    44
    5 == Agenda ==
     5== Notes ==
    66
    771. //[wiki:Research/LhARA/RadiationBiology/Meetings/2025-03-13 Minutes and actions]: //
    88 - **RR/KL:** Liaise with Will Shields and send to HO request for format of field map, e.g. B on grid with 0.5 mm pitch as starting point.
     9  - **Done.**
    910 - **JMcG/KL:** Get beam reference material onto wiki.
     11  - **Stands.**
    1012 - **CH/RG:** Seek to expedite procurement of stands etc.
     13  - **Done.** Design done.  Manufacture of quad holders is now ongoing.  Motorised stands ordered; waiting for them to be delivered.
    1114 - **EM/CD:** For next agenda: review Trello, add discussion of ideal beam time.
    12  - **KL, JMcG:** Collecting documentation for simulation etc. on wiki.
     15  - **Stands.**
    1316
    14172. //Field mapping, preparation of PMQs: // HO/BS \\
     18 - No report this time.
    1519
    16203. //Mechanical design update: // CH \\
     21 - Design complete.  CH will provide drawings to be placed in the repository/wiki.
    1722
    18234. //Simulation update: // AFR, CD \\
     24 - CD presented analysis of LION RCF stack data.  Noted there was a lot of low-intensity noise in the images.  Not sure of origin, but also, not clearly generated from the beam focusing.  Take a practical approach to the study.
    1925
    20265. //Bioprep: // EM/MH \\
     27 - MH: Rings have arrived and carossells are being made and are being anodised.
     28
     29 - EM by email: Two updates from my side, the glass rings have arrived and have been sent to Mark for final checks before the carousel is made. Also, frustratingly the first 14 vs 28 MeV test at Birmingham was unsuccessful. There was no survival in any of the irradiated dishes, although we have this same issue with our helium irradiations on the mylar dishes, and in both cases control dishes formed colonies. This is telling us that something is wrong with the irradiation set up with the rings – likely we think that they are drying out during the irradiation process. I have set up some tests to check how much media we need to leave on for the time of irradiation (at the moment we remove it all) and I should hopefully have some answers next week. We have more beam time booked in for 12th June to retest the protons. Although this is frustrating, it highlights the need for more optimisations to be done before we can do SCAPA experiments.
    2130
    22316. DoNM
     
    2433
    25347. AoB \\
     35 - Rehearse dates ...
     36  - 21May mtg at SCAPA with JPar and EM; RG will circulate agenda and see if appropriate to schedule a discussion of the way forward.
     37  - 16Jun--28Jun: Quad set up run;
     38  - 21Jul--01Aug: Bio exposure run.
    2639
    2740=== Summary of actions ===
    2841
     42 - **JMcG/KL:** Get beam reference material onto wiki.
     43 - **EM/CD:** For next agenda: review Trello, add discussion of ideal beam time.
     44 - **RG:*8 Circulate details of 21May when known.
     45
    2946
    3047----